OpenStudy (samigupta8):

if log 2 base x log 2 base 2x=log 2 base 4x then a.x=2^-√2 b.x=2^√2 c.x=1/4 d.x=2^2

OpenStudy (anonymous):

woW !

OpenStudy (anonymous):

so confusing, can;t you use the equation mode ? Do you mean :\[\log _{ x}2.\log _{ 2x}2 = \log _{ 4x}2\]

OpenStudy (samigupta8):

ryt....

OpenStudy (anonymous):

solve the equation (1/u)+(1/u+1)=(1/u+2)

OpenStudy (anonymous):

did you get sami?

OpenStudy (anonymous):

Do you know the formula : \[\log _{b}a = \frac{ \log _{n} a }{ \log _{n} b }\]

OpenStudy (samigupta8):

yaa

OpenStudy (anonymous):

use the property of logarithms log\[\log_a{b} =1/(\log_b{a}) \]

OpenStudy (anonymous):

if yes then 1. apply the formula both sides, 2. take log 2 = a, cause, its a constant anyway, & log x = y 3. simplify to get an algebraic equation in y 4. find y, then x

OpenStudy (anonymous):

ans is both A and B

OpenStudy (samigupta8):

how r u getting it

OpenStudy (anonymous):

i described the solution above

OpenStudy (samigupta8):

what is u here

OpenStudy (anonymous):

a variable equal to \[\log_2{x} \]

OpenStudy (samigupta8):

okk.i got it

OpenStudy (anonymous):

sure?